# OpenRLHF: An Easy-to-use, Scalable and High-performance RLHF Framework

### v2

@@ Abstract
As large language models (LLMs) continue to grow by scaling laws, reinforcement learning from human feedback (RLHF) has gained significant attention due to its outstanding performance. However, unlike pretraining or fine-tuning a single model, scaling reinforcement learning from human feedback (RLHF) for training large language models poses coordination challenges across four models. We present OpenRLHF, an open-source framework enabling efficient RLHF scaling. Unlike existing RLHF frameworks that co-locate four models on the same GPUs, OpenRLHF re-designs scheduling for the models beyond 70B parameters using Ray, vLLM, and DeepSpeed, leveraging improved resource utilization and diverse training approaches. Integrating seamlessly with Hugging Face, OpenRLHF provides an out-of-the-box solution with optimized algorithms and launch scripts, which ensures user-friendliness. OpenRLHF implements RLHF, DPO, rejection sampling, and other alignment techniques. Empowering state-of-the-art LLM development, OpenRLHF’s code is available at https://github.com/OpenLLMAI/OpenRLHF.
@@ 3.4 Ease of Use
1ray job submit -- python3 examples/train_ppo_ray.py \
@@
Listing 1: PPO startup method based on deepspeed (with Ray)
We provide usage scripts for all supported algorithms in the Appendix A.
@@ 5 Conclusion
Appendix A Appendix: Supported Algorithms And Usage
A.1 Preliminary Preparation
A.1.1 Supervised FineTuning
@@
1deepspeed train_sft.py \
2 --pretrain {Pre-trained model name or path} \
3 --dataset {SFT dataset name} \
4 --input_key {SFT dataset input key} \
5 --output_key {SFT dataset response key} \
6 --zero_stage 2 \
7 --bf16 \
8 --flash_attn \
9 --learning_rate 5e-6 \
10 --save_path {Model output path}
Listing 2: SFT startup method based on deepspeed
A.1.2 Reward Model Training
@@
1deepspeed train_rm.py \
2 --pretrain {Model name or path after SFT} \
3 --bf16 \
4 --zero_stage 3 \
5 --learning_rate 9e-6 \
6 --dataset {Pairwise preference dataset name} \
7 --chosen_key {Pairwise dataset chosen key} \
8 --rejected_key {Pairwise dataset rejected key} \
9 --flash_attn \
10 --save_path {Model output path}
Listing 3: RM startup method based on deepspeed
A.2 Proximal Policy Optimization
@@
A.3 Direct Preference Optimization
@@
1deepspeed train_dpo.py \
2 --train_batch_size 128 \
3 --micro_train_batch_size 1 \
4 --pretrain {Model name or path after SFT} \
5 --bf16 \
6 --zero_stage 3 \
7 --beta 0.1 \
8 --dataset {Pairwise preference dataset name} \
9 --chosen_key {Pairwise dataset chosen key} \
10 --rejected_key {Pairwise dataset rejected key} \
11 --flash_attn \
12 --ref_offload \
13 --save_path {Model output path}
Listing 4: DPO startup method based on deepspeed
A.4 Kahneman-Tversky Optimization
@@
1deepspeed train_kto.py \
2 --pretrain {Model name or path after SFT} \
3 --bf16 \
4 --zero_stage 3 \
5 --beta 0.1 \
6 --learning_rate 5e-7 \
7 --dataset {(Non-Pairwise or Pairwise) Preference dataset name} \
8 --prompt_key {Non-Pairwise dataset prompt key} \
9 --output_key {Non-Pairwise dataset output key} \
10 --label_key {Non-Pairwise dataset label key} \
11 --flash_attn \
12 --vanilla_loss \
13 --save_path {Model output path}
Listing 5: KTO startup method based on deepspeed
A.5 Rejection Sampling Finetuning
@@
1python batch_inference.py
2 --eval_task generate_vllm \
3 --pretrain {Model name or path after SFT} \
4 --dataset {Prompt dataset name} \
5 --intpu_key {Prompt dataset input key}
6 --temperature 0.9
7 --tp_size 8
8 --best_of_n 16 \
9 --output_path {GENERATED_SAMPLES}
Listing 6: RS GENERATE step startup method based on vLLM
1deepspeed batch_inference.py
2 --eval_task rm \
3 --pretrain {Reward model name or path} \
4 --bf16 \
5 --dataset {GENERATED_SAMPLES} \
6 --zero_stage 0 \
7 --post_processor rs \
8 --output_path {RS_FINETUNING_SAMPLES}
Listing 7: RS Score step startup method based on deepspeed
A.6 Conditional Supervised Finetuning

### v3

@@ Abstract
As large language models (LLMs) continue to grow by scaling laws, reinforcement learning from human feedback (RLHF) has gained significant attention due to its outstanding performance. However, unlike pretraining or fine-tuning a single model, scaling reinforcement learning from human feedback (RLHF) for training large language models poses coordination challenges across four models. We present OpenRLHF, an open-source framework enabling efficient RLHF scaling. Unlike existing RLHF frameworks that co-locate four models on the same GPUs, OpenRLHF re-designs scheduling for the models beyond 70B parameters using Ray, vLLM, and DeepSpeed, leveraging improved resource utilization and diverse training approaches. Integrating seamlessly with Hugging Face, OpenRLHF provides an out-of-the-box solution with optimized algorithms and launch scripts, which ensures user-friendliness. OpenRLHF implements RLHF, DPO, rejection sampling, and other alignment techniques. Empowering state-of-the-art LLM development, OpenRLHF’s code is available at https://github.com/OpenRLHF/OpenRLHF.
@@ 3.3 PPO Implementation Tricks
  • Initialize the Critic with the weights of the reward model.
  • Use a lower learning rate for the Actor while the Critic has a higher learning rate.
  • Freeze the weights of the Actor in the initial learning stage for better initialization of the Critic.
  • Use GAE (Generalized Advantage Estimation).
More details are in the blog [25].
@@ 3.4 Ease of Use
1pip install openrlhf[vllm]
2
3ray start --head --node-ip-address 0.0.0.0
4ray job submit -- python3 openrlhf.cli.train_ppo_ray \
@@
21 --apply_chat_template \ # Apply HF tokenizer template
@@
Listing 1: PPO startup method based on Deepspeed and Ray
We provide usage scripts for all supported algorithms (Appendix A) in the https://openrlhf.readthedocs.io/en/latest/usage.html.
@@ 4.2 Training Stability and Convergence
We provided the pre-trained checkpoints in https://huggingface.co/OpenRLHF.
@@ 5 Conclusion
Appendix A Appendix: Supported Algorithms
OpenRLHF currently supports the following alignment algorithms, with more in continuous development:
A.1 Supervised FineTuning
@@
A.2 Reward Model Training
@@
A.3 Proximal Policy Optimization
@@
A.4 Direct Preference Optimization
@@
A.5 Kahneman-Tversky Optimization
@@
A.6 Iterative Direct Preference Optimization
Iterative Direct Preference Optimization (Iterative DPO) [6] enhances vanilla DPO by continuously using an online approach to refine language models based on human feedback. The process involves three steps: (1) initially fine-tuning the model on instruction-following data, (2) collecting new preference data by sampling prompts and generating responses, and (3) iteratively updating the model with this new data. This method allows the model to adapt, improving its ability to generate high-quality responses and handle diverse inputs effectively.
A.7 Rejection Sampling Finetuning
@@
A.8 Conditional Supervised Finetuning